OpenStudy (anonymous):

an investor invested a total of 3,100 in two mutual funds. One fund earned a 6% profit while the other earned a 2% profit. If the investors total profit was $110 how much was invested in each mutual year.

OpenStudy (kropot72):

Let the amounts invested be A and B. The following equations can be set up from the given information: A + B = 3100 ........(1) 0.06A + 0.02B = 110 .........(2) We can find B in terms of A from equation (1) by rearrangement: B = 3100 - A .............(3) If you plug the expression for B given by (3) into equation (2) you can get an equation that can be solved to find the value of A. Can you do that?

OpenStudy (anonymous):

so it would be 3100*0.02+110

OpenStudy (kropot72):

0.06A + 0.02(3100 - A) = 110 0.06A + 0.02 * 3100 - 0.02A = 110 ..........(4)

OpenStudy (anonymous):

so I can get 2 separate answers wouldn't it be 0.06*3100 but where would the 110 fit in?

OpenStudy (kropot72):

There is only one answer for A: 0.06A + 0.02 * 3100 - 0.02A = 110 ..........(4) Gathering like terms in (4) gives: 0.06A - 0.02A + 0.02 * 3100 = 110 ........(5) Simplifying (5) gives: 0.04A + 62 = 110 .............(6) Now if you subtract 62 from the left hand side of (6) and subtract 62 from the right hand side of (6) what do you get?

OpenStudy (anonymous):

the way the answers are set up it wants to know the profit of 6% and then the profit of 2% underneath

OpenStudy (kropot72):

Amount A earns 6% and amount B earns 2%. You can see that from equation (2) in my first post. So now can you do the following to enable amount A to be found? 0.04A + 62 = 110 .............(6) If you subtract 62 from the left hand side of (6) and subtract 62 from the right hand side of (6) what do you get?

OpenStudy (kropot72):

After amount A is found, amount B can be found by pugging amount A into equation (1).

OpenStudy (kropot72):

The question that you have posted asks for the amounts invested, which are A and B.

OpenStudy (anonymous):

so a=48 im confused now

OpenStudy (kropot72):

No. Your value for A is incorrect. 0.04A + 62 = 110 .............(6) Subtracting 62 from both sides of (6) gives 0.04A + 62 - 62 = 110 - 62 ..........(7) Simplifying (7) gives 0.04A = 48 ............(8) Now divide both sides of (8) by 0.04 to find the value of A.

OpenStudy (kropot72):

0.04A = 48 ............(8) \[\frac{0.04A}{0.04}=A=\frac{48}{0.04}=you\ can\ calculate\]

OpenStudy (anonymous):

a=1200 b=

OpenStudy (kropot72):

Correct! Amount A that earns 6% is $1200. Now plug that value for A into equation (1) to find the amount B: A + B = 3100 ........(1) 1200 + B = 3100 ...........(9) Now subtract 1200 from both sides of (9).

OpenStudy (anonymous):

b=1900

OpenStudy (kropot72):

Correct! Amount A earning 6% is $1200 and amount B earning 2% is $1900.

OpenStudy (anonymous):

thx

OpenStudy (kropot72):

You're welcome :)